# Fixation and extinction in time-fluctuating spatially structured metapopulations

### v1

@@ A.2 Eco-evolutionary dynamics of a single deme
Since the metapopulation consists of a graph of connected demes, all with the same carrying capacity, we can gain significant insight into its dynamics by looking into its building block. In this section, we therefore analyse the eco-evolutionary dynamics of single isolated deme (when $m=0$).
For an isolated deme, there is only intra-deme dynamics according to the birth-death process defined by Eqs. (1) and (2). The ME for the dynamics of a single isolated deme is thus given by setting $\Omega=1$ and $m=0$ in Eq. (27), see Appendix A.1. The corresponding intra-deme dynamics in a time-varying environment can thus be simulated using exact methods like the Gillespie algorithm [98], as in Refs [16, 17, 18, 23, 20, 21, 22, 24]. It is instructive to ignore all forms of fluctuations and consider the mean-field approximation of an isolated deme dynamics subject to a constant carrying capacity $K\gg 1$. Following Refs. [16, 17], with the transition rates Eq. (2), the mean-field eco-evolutionary dynamics of a single isolated deme is characterised by rate equations for the size $n$ of the deme and the fraction $x=n_{M}/n$ of mutants in the deme, which read:
@@
where the dot indicates the time derivative and we have used $f_{M}=1+s$ and $f_{W}=1$. These decoupled rate equations predict the quick relaxation of the deme size towards the constant carrying capacity, with $n\to K$ on a timescale $t\sim 1$, and the growth of the fraction of mutants, with $x\to 1$ on a timescale $t\sim 1/s$. In the total absence of fluctuations, when $0<s\ll 1$ (small selective advantage to $M$), in the mean-field picture, the deme size quickly approaches the carrying capacity and there is a timescale separation between $n$ and $x$, respectively the fast and slow variables. $W$ cells are thus slowly wiped out by mutants that take over the deme on a timescale $t\sim 1/s\gg 1$ [16, 17]. It is worth noting that with the effective transition rates Eq. (19) we have assumed that invasions always occur after deme size and composition relaxation. This means that Eqs. (23) and (24) assume a timescale separation between $n$ and $x$ and $x$ and the number of $M$ demes.
@@
When demographic fluctuations can be neglected and randomness only arises from enviromental variability via Eq. (5), the intra-deme dynamics of an isolated deme is well-captured by the piecewise deterministic Markov process (PDMP) for $n$, obtained by ignoring demographic fluctuations [84, 16, 17]. In the realm of the PDMP approximation, the deme size thus satisfies a deterministic logistic equation in each environmental state, subject to a carrying capacity that switches when the environment changes ($K=K_{\pm}$ when $\xi=\pm 1$), yielding the following $n$-PDMP [16, 17, 18, 22, 32, 21]:
@@
$p(\nu,n)=\frac{1}{2}\sum_{\xi}p_{\xi}(\nu,n)=\frac{\mathcal{Z}}{n^{2}}\left[\left(\frac{K_{+}}{n}-1\right)\left(1-\frac{K_{-}}{n}\right)\right]^{\nu-1},$ (30)
@@
where $\mathcal{Z}$ is the normalisation constant and $n\in[K_{-},K_{+}]$. Despite ignoring the effect of demographic noise, $p(\nu,n)$ aptly captures many properties of the quasi-stationary distribution of the size of an isolated single deme [17, 18]. For instance, the long-time average deme size is accurately approximated by ${\cal N}(\nu)=\int_{K_{-}}^{K_{+}}np(\nu,n){\rm d}n$, and is a decreasing function of $\nu$ [16, 17]. As illustrated by Fig. 7, the density $p(\nu,n)$ correctly predicts that the deme size distribution is bimodal when $\nu<1$ and unimodal when $\nu>1$, and that it is peaked at $n\approx K_{\pm}$ when $\nu\ll 1$ (slow switching) and centred around $n\approx{\cal K}$ when $\nu\gg 1$ (fast switching), see Eq. (15). The joint and marginal PDMP probability densities Eq. (17) and Eq. (30) provide valuable insight into the deme size distribution when these are subject to weak bottlenecks and their extinction can be neglected [16, 17, 18, 22, 32, 21, 19].
While $p_{\xi}(\nu,n)$ and $p(\nu,n)$ give a PDMP description of the quasi-stationary distribution of the size of an isolated deme ($m=0$), the $n$-PDMP stationary densities (17) and (30) are still a valid approximations of the long-time size distribution of $n$ in the presence of migration as considered here. In fact, as shown below in Fig. 7, the influence of migration on the distribution of the deme size is essentially unnoticeable, and its main features are therefore well captured by Eq. (17) and Eq. (30).
@@
Figure 7: Quasi-stationary probability density of deme population size ($n$-QPD) distribution for various parameters and its $n$-PDMP approximation by (17) and (30). Red and blue bars show data for the $n$-QPD conditioned on $K(t)=K_{-}$ and $K(t)=K_{+}$, respectively. Red, blue, and black solid lines are the $n$-PDMP stationary densities $p_{-}(\nu,n)$, $p_{+}(\nu,n)$, and $p(\nu,n)$, respectively, given by Eqs. (17) and (30). (a-c) are for clique, (d-f) for cycle, and (g-i) for grid metapopulations. We have $\nu=10^{-3}$ in (a,d,g), $\nu=10^{-1}$ in (b,e,h), and $\nu=10^{2}$ in (c,f,i). Other parameters are $\Omega=16$, $m=10^{-3}$, $K_{+}=200$, and $K_{-}=20$. All represent a single realisation tracked until $t=10^{5}$.
@@
In the main text, we have used the PDMP approximation to describe the size distribution of a single isolated deme subject to symmetric random switching of its carrying capacity Eq. (5). Here, we show that spatial migration has no noticeable influence on the size distribution of a single deme of a metapopulation structured as a regular circulation graph ${\rm G}=\{{\rm clique,cycle,grid}\}$. To this end, in Fig. 7, we compare the size distribution of a single deme in a metapopulation structured as a regular ${\rm G}$-graph in the presence of a migration per capita rate $m=10^{-3}$ (obtained from stochastic simulations) with the predictions of Eq. (17). These results illustrate that migration has no noticeable effect on the deme size distribution that can approximated by PDMP density Eq. (17) (or Eq. (30)) in the absence and presence of migration on any graph ${\rm G}=\{{\rm clique,cycle,grid}\}$
@@
As a consequence, the deme size distribution of any metapopulation considered here is well approximated by the joint and marginal PDMP densities $p(\pm,\nu,n)$ and $p(\nu,n)$, given by Eq. (17) and Eq. (30).
@@
Intuitively, this can be understood by noticing that the spatial structures considered here are circulation graphs, yielding the same inward and outward migration flow at each deme, and each deme has the same carrying capacity. As a consequence, the average number of cells per deme is expected to be independent of migration. The latter remains well captured by $p(\pm,\nu,n)$ and $p(\nu,n)$ on all spatial structures considered here, as seen in Fig. 7.
Appendix B Deme invasion and slow migration
@@
In this section, we analyse the process of invasion of a single deme subject to a constant carrying capacity $K$ when $\psi(m,K)\gg 1$, see Sec. III.0.1. In this competition-dominated regime, the extinction of demes can be neglected and their size rapidly fluctuated about $K$, see Fig. 1(a). In this scenario, we can assume that the deme size is constant $n(x)=K$, and describe the deme dynamics by tracking the number of mutants $n_{M}$ and wild-type cells $n_{W}$ in the deme $x$. The deme composition $(n_{M},n_{W})=(n_{M},K-n_{M})$ thus changes according to the Moran process [64, 7, 65, 66]
@@
These transition rates correspond to the the effective rates of increase and decrease in the number of $M$ in a deme of size $K$. This Moran process conserves the deme size by accompanying each birth of an $M/W$ by the simultaneous death of a $W/M$, and is characterised by the absorbing states $(n_{M},n_{W})=(K,0)$ ($M$ deme) and $(n_{M},n_{W})=(K,0)$ ($W$ deme). The $M$ fixation probability $\phi_{\text{Mo}}$ and unconditional mean fixation time (uMFT) $\theta_{\text{Mo}}$ for this Moran process are classical results, and when there are initially $i$ cells of type $M$, they read [7, 67, 65, 66]
$\phi_{\text{Mo}}(i)$ $=\frac{1-\gamma_{\text{Mo}}^{i}}{1-\gamma_{\text{Mo}}^{K}},$ (33)
$\theta_{\text{Mo}}(i)$ $=\phi_{\text{Mo}}(i)\sum_{n=i}^{K-1}\sum_{l=1}^{n}\frac{\gamma_{\text{Mo}}^{n-l}}{{\cal T}_{\text{Mo}}^{+}(l)}$
$-(1-\phi_{\text{Mo}}(i))\sum_{n=1}^{i-1}\sum_{l=1}^{n}\frac{\gamma_{\text{Mo}}^{n-l}}{{\cal T}_{\text{Mo}}^{+}(l)},$
where $\gamma_{\text{Mo}}\equiv{\cal T}_{\text{Mo}}^{-}/{\cal T}_{\text{Mo}}^{+}=f_{W}/f_{M}=1/(1+s)$. The fixation probability of a single mutant ($i=1$) and of a single $W$ cell ($i=K-1$) are particularly relevant for our purposes, and explictily read
@@
In the competition-dominated regime, $1/s$ sets the timescale of deme dynamics, see Eqs. (28). The growth of the cluster of $i$ mutant demes in time $1/s$ is $T^{+}(m,{\rm G},K)/s=mKE_{{\rm G}}(i)\rho_{M}/(q_{{\rm G}}s)$. In the adopted coarse-grained description, slow migration is the regime where the invasion can be regarded as being instantaneous, with fixation of a successful $M$ invader occurring before the next invasion. This requires that the average number of successful $M$ invaders in the time for the on-deme dynamics is less than one, $T^{+}(m,{\rm G},K)/s<1$, leading to the following condition for slow migration:
$m<\frac{sq_{{\rm G}}}{K\rho_{M}E_{{\rm G}}(i)}<\frac{s}{K\rho_{M}},$ (35)
where we have used $q_{\rm G}/E_{\rm G}(i)\leq 1$. When $s\ll 1$ and $Ks\gg 1$, we have $\rho_{M}\sim s$ and therefore simply $m<1/K$. For $\Omega=16,K=100$ and $s=0.1$, we can estimate that there is slow migration if $m\lesssim 10^{-2}$, which is in line with the values of $m\in[10^{-5},10^{-2}]$ used in our examples, and matches where we see deviations in Fig. 2(d).
@@ Appendix C Deme and metapopulation mean extinction times
Figure 8: $\theta_{E}$ vs. metapopulation size $\Omega$ for $K=7$ (yellow), $K=5$ (blue), $K=3$ (red) and $m=10^{-4}$. Markers are simulation results and lines are predictions of Eq. (39) for cliques (solid lines / crosses), cycles (dashed lines / circles), and grids (dotted lines / triangles). Markers of the same colour are almost indistinguishable. Deviations occur due to the approximation of $\tau_{E}(K)$ in Eq. (37). Selection plays no role in this regime, so results have been obtained with $s=0$.
In this section, we discuss the process of extinction of a single deme that has a constant carrying capacity $K$ when $\psi(m,K)\ll 1$, see Sec. III.0.2. In this extinction-dominated regime, we can assume that the deme size rapidly fluctuates about $K$, and extinction occurs from a deme of constant size $n(x)=K$ prior to any invasion. Without loss of generality (see below), we hence assume that extinction occurs from entirely occupied demes, with the metapopulation consisting only of $W$ and $M$ demes, all of size $K$. In this representation, the dynamics of a $W/M$ deme when $\psi\ll 1$ is given by the birth-death process of Sec. II with transition rates $T_{W/M}^{+}=n_{W/M}$ and $T_{W/M}^{-}=n_{W/M}^{2}/K$, subject to an absorbing boundary at $n_{W/M}=0$, see Eq. (2). Clearly therefore, the deme dynamics is independent of its type, and the deme mean extinction time (dMET) is the average time to reach $n_{W/M}=0$ and is the same for $W$ and $M$ demes (dMET is independent of $s$). A classical calculation yields (see, e.g. Sec. 6.7 in Ref. [71]) yields the following expression for an initially fully occupied deme of size $K$:
@@
The metapopulation mean extinction time (mMET) in the regime $\psi(m,K)\ll 1$ can be obtained analytically within the realm of the above coarse-grained description, in the spirit of the approach of Ref. [59] for cliques. The metapopulation thus consists initially of entirely occupied demes ($i$ mutant demes and $\Omega-i$ type $W$ demes). Since deme extinction here occurs prior to any invasion, we describe the metapopulation dynamics in terms of the number $j=0,1,\dots,\Omega$ of entirely occupied demes. Residents (of either $W$ or $M$ type) of these filled demes can recolonise a neighbouring empty site at a rate $B(j)$, see Fig. 1(d). In addition, each occupied deme goes extinct at a rate $D(j)$. This coarse-grained description of the metapopulation dynamics is therefore a birth-death process with an absorbing state $j=0$ corresponding to the eventual extinction of the metapopulation, and a reflecting boundary at $j=\Omega$ (all demes are occupied). In this picture, proceeding as above [71], the mMET reads
@@
In the vein of Ref. [59] the recolonisation-birth rate of occupied demes is $B(j)=mKj(1-j/\Omega)$, corresponding to a logistic growth with a rate proportional to the the expected number of migrations from an occupied deme $mK$. Here, the extinction rate is $D(j)=j/\tau_{E}$ and is inversely proportional to the mean local extinction time, that is the dMET. With $\psi=mK\tau_{E}$, using $\prod_{l=n}^{j-1}(1-\frac{l}{\Omega})=\frac{1}{\Omega^{j-n}}\frac{(\Omega-n)!}{(\Omega-j)!}$, Eq. (38) can be rewritten as
@@
The result $\theta_{E}(K,\Omega)$ has been explicitly derived for cliques (island model), but is found to provide good qualitative insight into the extinction dynamics for cycles and grids, see Fig. 2

### v2

@@ A.2 Eco-evolutionary dynamics of a single deme
Since the metapopulation consists of a graph of connected demes, all with the same carrying capacity, we can gain significant insight into its dynamics by looking into its building block. In this section, we therefore analyse the eco-evolutionary dynamics of a single isolated deme (when $m=0$).
For an isolated deme, there is only intra-deme dynamics according to the birth-death process defined by Eqs. (1) and (2). The ME for the dynamics of a single isolated deme is thus given by setting $\Omega=1$ and $m=0$ in Eq. (28); see Appendix A.1. The corresponding intra-deme dynamics in a time-varying environment can be simulated using exact methods like the Gillespie algorithm [97], as in Refs [16, 17, 18, 23, 20, 21, 22, 24]. It is instructive to ignore all forms of fluctuations and consider the mean-field approximation of an isolated deme dynamics subject to a constant carrying capacity $K\gg 1$. Following Refs. [16, 17], with the transition rates Eq. (2), the mean-field eco-evolutionary dynamics of a single isolated deme is characterised by rate equations for the size $n$ of the deme and the fraction $x=n_{M}/n$ of mutants in the deme, which read:
@@
where the dot indicates the time derivative and we have used $f_{M}=1+s$ and $f_{W}=1$. The prefactors of these decoupled rate equations predict the relaxation of the deme size towards the constant carrying capacity, with $n\to K$ on a timescale $t\sim 1$, and the growth of the fraction of mutants, with $x\to 1$ on a timescale $t\sim 1/s$. In the total absence of fluctuations, when $0<s\ll 1$ (small selective advantage to $M$), in the mean-field picture, the deme size quickly approaches the carrying capacity and there is a timescale separation between $n$ and $x$, respectively the fast and slow variables, where $n$ evolves on a timescale of order $1$ and $x$ on a timescale $\sim 1/s\gg 1$; see Table 1. Thus, $W$ cells are thus slowly wiped out by mutants that take over the deme on a timescale $t\sim 1/s\gg 1$ [16, 17]. It is worth noting that with the effective transition rates Eq. (21) we have assumed that invasions always occur after deme size and composition relaxation. This means that Eqs. (25) and (26) assume a timescale separation between $n$ and $x$ in addition to the slow migration assumption.
@@
When demographic fluctuations can be neglected and randomness only arises from environmental variability via Eq. (6), the intra-deme dynamics of an isolated deme is well-captured by the piecewise deterministic Markov process (PDMP) for $n$, obtained by ignoring demographic fluctuations [85, 16, 17]. In the realm of the PDMP approximation, the deme size thus satisfies a deterministic logistic equation in each environmental state, subject to a carrying capacity that switches when the environment changes ($K=K_{\pm}$ when $\xi=\pm 1$), yielding the following $n$-PDMP [16, 17, 18, 22, 32, 21]:
@@
$p(n;\nu)=\frac{1}{2}\sum_{\xi}p_{\xi}(n;\nu)=\frac{\mathcal{Z}}{n^{2}}\left[\left(\frac{K_{+}}{n}-1\right)\left(1-\frac{K_{-}}{n}\right)\right]^{\nu-1},$ (31)
@@
where $\mathcal{Z}$ is the normalisation constant and $n\in[K_{-},K_{+}]$. Despite ignoring the effect of demographic noise, $p(n;\nu)$ aptly captures many properties of the quasi-stationary distribution of the size of an isolated single deme [17, 18]. For instance, the long-time average deme size is accurately approximated by ${\cal N}(\nu)=\int_{K_{-}}^{K_{+}}np(n;\nu){\rm d}n$, and is a decreasing function of $\nu$ [16, 17].
While $p_{\xi}(n;\nu)$ and $p(n;\nu)$ give a PDMP description of the quasi-stationary distribution of the size of an isolated deme ($m=0$), the $n$-PDMP stationary densities given by Eqs. (19) and (31) are still a valid approximation of the long-time size distribution of $n$ in the presence of migration between many connected demes, as considered here. In fact, as shown below in Fig. 7, the influence of migration on the distribution of the deme size is essentially unnoticeable, where the predictions of Eqs. (19) for a single isolated deme (no migration) are compared with the deme size distribution obtained from by sample averaging across all demes of the metapopulation in the presence of migration. Its main features are therefore well captured by Eq. (19) and Eq. (31). As illustrated by Fig. 7, the density $p(n;\nu)$ correctly predicts that the deme size distribution is bimodal when $\nu<1$ and unimodal when $\nu>1$, and that it is peaked at $n\approx K_{\pm}$ when $\nu\ll 1$ (slow switching) and centred around $n\approx{\cal K}$ when $\nu\gg 1$ (fast switching), see Eq. (17). The joint and marginal PDMP probability densities Eq. (19) and Eq. (31) provide valuable insight into the deme size distribution when these are subject to weak bottlenecks and their extinction can be neglected [16, 17, 18, 22, 32, 21, 19].
@@
Figure 7: Quasi-stationary probability density of deme population size ($n$-QPSD) distribution on clique metapopulations for various parameters and its $n$-PDMP approximation given by Eqs. (19) and (31). Red and blue bars show data for the $n$-QPSD conditioned on $K(t)=K_{-}$ and $K(t)=K_{+}$, respectively. Red, blue, and black solid lines are the $n$-PDMP stationary densities $p_{-}(n;\nu)$, $p_{+}(n;\nu)$, and $p(n;\nu)$, respectively, given by Eqs. (19) and (31). (a-c) are for $m=10^{-3}$, (d-f) for $m=10^{-2}$, and (g-i) for $m=10^{-1}$. We have $\nu=10^{-3}$ in (a,d,g), $\nu=10^{-1}$ in (b,e,h), and $\nu=10^{2}$ in (c,f,i). Other parameters are $\Omega=16$, $K_{+}=200$, and $K_{-}=20$. All represent a single realisation tracked until $t=10^{5}$.
@@
In the main text, we have used the PDMP approximation to describe the size distribution of a single isolated deme subject to symmetric random switching of its carrying capacity Eq. (6). Here, we show that spatial migration has no noticeable influence on the size distribution of a single deme of a metapopulation structured as a clique (though the same holds for other spatial structures). To this end, in Fig. 7, we compare the size distribution of a single deme in a metapopulation structured as a clique in the presence of a migration per capita rate $m\in\{10^{-3},10^{-2},10^{-1}\}$ (obtained from stochastic simulations) with the predictions of Eq. (19). These results illustrate that migration has no noticeable effect on the deme size distribution that can approximated by PDMP density Eq. (19) (or Eq. (31)) in the absence and presence of migration.
@@
As a consequence, the deme size distribution of any metapopulation considered here is well approximated by the joint and marginal PDMP densities $p_{\pm}(n;\nu)$ and $p(n;\nu)$, given by Eq. (19) and Eq. (31).
@@
Intuitively, this can be understood by noticing that the spatial structures considered here are circulation graphs, yielding the same inward and outward migration flow at each deme, and each deme has the same carrying capacity. As a consequence, the average number of cells per deme is expected to be independent of migration. The latter remains well captured by $p_{\pm}(n;\nu)$ and $p(n;\nu)$ regardless of migration rate, as seen in Fig. 7.
Appendix B Deme invasion
@@
In this section, we analyse the process of invasion of a single deme subject to a constant carrying capacity $K$ when $\psi(m,K)\gg 1$; see Sec. III.1. In this competition-dominated regime, the extinction of demes can be neglected and their size rapidly fluctuates about $K$; see Fig. 1(a). In this scenario, we can assume that the deme size is constant $n(x)=K$, and describe the deme dynamics by tracking the number of mutants $n_{M}$ and wild-type cells $n_{W}$ in the deme $x$. The deme composition $(n_{M},n_{W})=(n_{M},K-n_{M})$ thus changes according to the Moran process [64, 7, 65, 66]
@@
These transition rates correspond to the the effective rates of increase and decrease in the number of $M$ in a deme of size $K$. This Moran process conserves the deme size by accompanying each birth of an $M/W$ by the simultaneous death of a $W/M$, and is characterised by the absorbing states $(n_{M},n_{W})=(K,0)$ ($M$ deme) and $(n_{M},n_{W})=(0,K)$ ($W$ deme). The $M$ fixation probability $\phi_{\text{Mo}}$ and unconditional mean fixation time (uMFT) $\theta_{\text{Mo}}$ for this Moran process are classical results, and when there are initially $i$ cells of type $M$, they read [7, 67, 65, 66]
$\hskip-8.53581pt\begin{aligned} \phi_{\text{Mo}}(i)&=\frac{1-\gamma_{\text{Mo}}^{i}}{1-\gamma_{\text{Mo}}^{K}},\\ \theta_{\text{Mo}}(i)&=\phi_{\text{Mo}}(i)\sum_{n=i}^{K-1}\sum_{l=1}^{n}\frac{\gamma_{\text{Mo}}^{n-l}}{{\cal T}_{\text{Mo}}^{+}(l)}\\ &-(1-\phi_{\text{Mo}}(i))\sum_{n=1}^{i-1}\sum_{l=1}^{n}\frac{\gamma_{\text{Mo}}^{n-l}}{{\cal T}_{\text{Mo}}^{+}(l)},\end{aligned}$ (34)
where $\gamma_{\text{Mo}}\equiv{\cal T}_{\text{Mo}}^{-}/{\cal T}_{\text{Mo}}^{+}=f_{W}/f_{M}=1/(1+s)$. The fixation probability of a single mutant ($i=1$) and of a single $W$ cell (i=K1i=K-1) are particularly relevant for our purposes, and explicitly read
@@ Appendix C Deme and metapopulation mean extinction times
Figure 8: $\theta_{E}$ vs. metapopulation size $\Omega$ for $K=7$ (yellow), $K=5$ (blue), $K=3$ (red) and $m=10^{-4}$. Markers are simulation results and lines are predictions of Eq. (39) for cliques (solid lines / crosses), cycles (dashed lines / circles), and grids (dotted lines / triangles). For the grid, the values of $\Omega$ are chosen to maintain a square lattice, and thus occasionally differ slightly from the metapopulation size used in the cycle and clique. Markers of the same colour are almost indistinguishable. Deviations occur due to the approximation of $\tau_{E}(K)$ in Eq. (37). Selection plays no role in this regime, so results have been obtained with $s=0$.
In this section, we discuss the process of extinction of a single deme that has a constant carrying capacity $K$ when $\psi(m,K)\ll 1$; see Sec. III.2. In this extinction-dominated regime, we can assume that the deme size rapidly fluctuates about $K$, and extinction occurs from a deme of constant size $n(x)=K$ prior to any invasion. Without loss of generality (see below), we hence assume that extinction occurs from entirely occupied demes, with the metapopulation consisting only of $W$ and $M$ demes, all of size $K$. In this representation, the dynamics of a $W/M$ deme when $\psi\ll 1$ is given by the birth-death process of Sec. II with transition rates $T_{W/M}^{+}=n_{W/M}$ and $T_{W/M}^{-}=n_{W/M}^{2}/K$, subject to an absorbing boundary at $n_{W/M}=0$; see Eq. (2). Clearly therefore, the deme dynamics is independent of its type, and the deme mean extinction time (dMET) is the average time to reach $n_{W/M}=0$ and is the same for $W$ and $M$ demes (dMET is independent of $s$). A classical calculation (see, e.g. Sec. 6.7 in Ref. [72]) yields the following expression for an initially fully occupied deme of size $K$:
@@
The metapopulation mean extinction time (mMET) in the regime $\psi(m,K)\ll 1$ can be obtained analytically within the realm of the above coarse-grained description, in the spirit of the approach of Ref. [59] for cliques. The metapopulation thus consists initially of entirely occupied demes ($i$ mutant demes and $\Omega-i$ type $W$ demes). Since deme extinction here occurs prior to any invasion, we describe the metapopulation dynamics in terms of the number $j=0,1,\dots,\Omega$ of entirely occupied demes. Residents (of either $W$ or $M$ type) of these filled demes can recolonise a neighbouring empty site at a rate $B(j)$; see Fig. 1(d). In addition, each occupied deme goes extinct at a rate $D(j)$. This coarse-grained description of the metapopulation dynamics is therefore a birth-death process with an absorbing state $j=0$ corresponding to the eventual extinction of the metapopulation, and a reflecting boundary at $j=\Omega$ (all demes are occupied). In this picture, proceeding as above [72], the mMET reads
@@
In the vein of Ref. [59] the recolonisation-birth rate of occupied demes is $B(j)=mKj(1-j/\Omega)$, corresponding to a logistic growth with a rate proportional to the expected number of migrations from an occupied deme $mK$. Here, the extinction rate is $D(j)=j/\tau_{E}$ and is inversely proportional to the mean local extinction time, that is the dMET. With $\psi=mK\tau_{E}$, using $\prod_{l=n}^{j-1}(1-\frac{l}{\Omega})=\frac{1}{\Omega^{j-n}}\frac{(\Omega-n)!}{(\Omega-j)!}$, Eq. (38) can be rewritten as
@@
The result $\theta_{E}(K,\Omega)$ has been explicitly derived for cliques (island model), but is found to provide good qualitative insight into the extinction dynamics for cycles and grids; see Fig. 2.